# LiRA: Light-Robust Adversary for Model-based Reinforcement Learning in Real World

### v1

@@ Abstract
journal: Neural Networks
url]http://tk.prinlab.org/en/
@@ 1 Introduction
Reinforcement learning (RL) attracts increasing attention as methodology for robots to learn stochastic control policies that enable them to accomplish a given task by trial and error (Sutton and Barto, 2018). In particular, model-based RL has high expectations for real-world robotic applications due to its excellent sample efficiency (Chua et al., 2018; Williams et al., 2018), and various robotic applications have been reported recently. : e.g. manipulation on complex dynamics of dexterous hand (Nagabandi et al., 2020), cloth (Luque et al., 2024), and even grinding of object (Hachimine et al., 2023); locomotion by quadruped (Yang et al., 2020), biped (Kuo et al., 2023), and quadrotor (Lambert et al., 2019); and their combination (Arcari et al., 2023). Although the framework proposed in this study can be applied to model-free RL, for the sake of simplicity, this paper limits its focus to model-based RL.
@@
In addition to the sample efficiency (and, of course, control performance), recent RL studies have often aim to maximize robustness. This would enable robots to cope with unexpected situations in a faced environment, which can easily be caused by a variety of factors (e.g. biases in empirical data and unobservable information). A naive approach is domain randomization on simulations (Tobin et al., 2017; Ramos et al., 2019; Muratore et al., 2022), which encompasses a variety of optimal policies by learning from data experienced in simulations driven by various simulation parameters. Although this approach has had much success in recent years, it should be inefficient unless data are collected in parallel using plenty of computational resources, which waste huge amounts of energy (Andrychowicz et al., 2020; Rudin et al., 2022). It is therefore difficult for real-world robots to continuously learn on-board with this approach.
As a more efficient approach, an adversarial learning can be considered, in which an adversary is introduced that actively interferes with the robot’s task accomplishment (Pinto et al., 2017; Gleave et al., 2020; Zhai et al., 2022). Such active disturbances can produce complex and challenging experiences more efficiently than random ones, but this approach inevitably inherits the mode collapse issue in adversarial learning (Srivastava et al., 2017; Liu et al., 2019), making learning the optimal policy unstable. In addition, in real-world robots, excessive disturbances might damage themselves and/or objects in the faced environment.
Furthermore, one of the adverse effects of maximizing robustness (caused in both of the above) is to make the policy too conservative (Petrik and Russel, 2019; Lechner et al., 2023; Huang et al., 2023). In other words, since robots have to take into account contingencies that may not occur in practice, they tend to choose safer and more secure actions. If the maximum disturbance intensity could appropriately be designed in advance, such conservativeness would be controlled at the minimum. However, as the relationship between the policy performance and the disturbance intensity is nonlinear and varies from situation to situation, this solution is infeasible.
This issue would be mitigated by domain adaptation (Yu et al., 2017; Semage et al., 2022; Ilboudo et al., 2023b), which estimates the uncertainty in the faced environment (e.g. represented as simulation parameters and disturbances) and incorporates them into the policy. As long as the uncertainty estimator is accurate, this approach can acquire the necessary and sufficient balance between robustness and conservativeness. However, its learning cost should be equal to or larger than that of domain randomization, since the effects of any potential uncertainty must be learned in advance.
In addition, the extra computational cost of the uncertainty estimator and the uncertainty-conditioned policy can easily be an obstacle to real-world robot control, which is strictly time-constrained. In particular, it has been reported that in the model-based RL algorithms, where model predictive control (MPC) (Williams et al., 2018; Kobayashi and Fukumoto, 2022) is employed in combination, reducing the extra computational cost as much as possible leads to better control performance because the MPC optimization converges in time (Lenz et al., 2015; Kobayashi and Watanuki, 2023). As the estimation error would also be raised as our concern, an alternative solution to this domain adaptation should be considered.
@@
In this context summarized in Table 1, this study proposes a new framework, so-called LiRA (see Fig. 1), for adversarial learning that satisfies the following two requirements:
@@
To this end, adversarial learning is first rederived according to variational inference (Kingma and Welling, 2014). Then, as a new definition of robustness, light robustness (Fischetti and Monaci, 2009) is integrated with it. This relaxes full robustness by imposing an inequality constraint that limits the degradation from ideal performance due to disturbances within a specified threshold (this corresponds to the conservativeness). Through the Lagrangian method (Kobayashi, 2023), this inequality constraint is transformed into a loss function for numerical optimization, allowing for automatic tuning of the adversary level. In addition, as such a learning process makes the model marginalized w.r.t. disturbances moderately robust, the disturbances can remain unknown at inference time, requiring no estimation cost or additional conditioning.
The behaviors of the proposed LiRA are confirmed by numerical simulations using Mujoco (Todorov et al., 2012). LiRA is not too conservative in control performance only with the nominal noise, and the degradation of control performance in response to disturbance intensity is suppressed. In addition, according to the specified conservativeness, the proposed LiRA attempts to increase the robustness by increasing the adversary level for the condition with weaker disturbance sensitivity, and vice versa. Such an auto-tuning capability also yields self-paced curriculum learning (Kumar et al., 2010; Klink et al., 2020), where disturbances are suppressed during the under-performance phase and reinforced as learning progresses. These results indicate that LiRA can achieve a good balance between robustness and conservativeness.
Finally, a real-robot learning using LiRA is demonstrated without simulations. By applying virtual external forces to its internal impedance model, a quadrupedal robot perceives disturbances only during learning. This task can be performed by LiRA only on the real robot with appropriate suppression of excessive disturbance, whereas a fully adversarial learning has a high risk of malfunctioning the robot. Although learning with nominal noise alone should not be sufficient to generate a force-responsive gait, LiRA makes it possible by moderately experiencing various situations during learning. Thus, this demonstration illustrates that LiRA can induce the robot to acquire the desired behaviors only with the limited data collected from the real world in short time (less than two hours in this case).
@@ 2.1 Model-based reinforcement learning
In RL (Sutton and Barto, 2018), an agent aims to gain the maximum rewards from an environment in the future. For mathematical formulation of the relationship between the agent and environment, Markov decision process (MDP) is basically assumed, i.e. $(\mathcal{S},\mathcal{A},p_{e},r)$. Here, $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $p_{e}:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$ is the state transition probability (a.k.a. dynamics), and $r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathcal{R}$ is the reward function to evaluate each transition. Here, $\mathcal{R}\subseteq\mathbb{R}$ holds normally, but the case of multi-objective RL, where multiple rewards are simultaneously considered (Hayes et al., 2022), is assumed for generality. That is, the output of reward function is vectorized as $\mathcal{R}\subseteq\ \mathbb{R}^{|\mathcal{R}|}$.
@@
If the agent knows $p_{e}$ and $r$ accurately, this problem can numerically be solved using MPC (in this paper, AccelMPPI (Kobayashi and Fukumoto, 2022) is employed). Therefore, model-based RL algorithms explicitly approximate them as a model ($p_{w}$ in this paper) using, for example, deep neural networks with parameters $\theta$ (Chua et al., 2018; Kobayashi and Watanuki, 2023).
@@ 2.2 Adversarial learning
To make $\pi^{\ast}$ obtained through the above optimization problem robust, worst-case scenarios for the objective function (i.e. the sum of predicted reward, called the return) are basically considered (Köhler et al., 2020; Zanon and Gros, 2020). However, this incurs an extra computational cost at inference time compared to the case simply using expected value of the prediction. In this study, therefore, robust control is achieved by making $p_{w}$ robust during learning, referring to the result reported in the literature (Aotani and Kobayashi, 2024) that control performance becomes implicitly robust if $p_{w}$ is optimized in consideration of events that are rare in reality.
@@ 3.1 Overview
In order to improve the robustness of the control performance to the extent that it does not become too conservative, this paper proposes LiRA. LiRA utilizes two theoretical tricks in its formulation. First, the adversarial learning in eq. (3) is reformulated by variational inference (Kingma and Welling, 2014), introducing the necessary components in a natural way. Then, the light robustness (Fischetti and Monaci, 2009) is applied as an additional constraint for the gap of predictive performance of models with/without disturbance information, which can be converted to a regularization term according to Lagrangian method with an auto-tuned weight parameter. This parameter has a role of adjusting adversary level automatically, balancing the robustness and conservativeness.
To appropriately solve the derived optimization problem, this paper implements three practical tricks (see the section 4). First, noting that the bias due to disturbances should ideally be zero, a design constraint is provided by a restricted normalizing flows (RNF) (Kobayashi and Aotani, 2023) such that the expected values of the two models are consistent with each other. Second, to make a gradient reversal layer (GRL) (Ganin et al., 2016) available for the stable and efficient adversarial learning, a trick to hindsight the computational graph from the disturbance generator to the model via the disturbance, so-called hindsight reparameterization gradient (HRG), is newly developed. Third, a new equality constraint to actively reflect the worst cases for either the model or the disturbance generator, midrange-mean balancing (MMB), is given, following the literature (Aotani and Kobayashi, 2024), which efficiently improves the robustness of the model by (moderately) prioritizing the worst-case data for the performance.
The details of the above proposal will be presented in the following sections. The pseudocode of LiRA is shown in Alg. 1. Note that the loss to be minimized in this pseudocode, $\mathcal{L}^{x}$, has the computational graph only for $x$ and cuts for the other variables to be optimized. In addition, $\mathcal{L}^{\phi}$ corresponds to Kullback-Leibler divergence approximated by Monte Carlo method, which is necessary as the adversary is designed by normalizing flows.
@@
19:         $\ell^{a}_{i}=-\ln p_{w}(s^{\prime}_{i},r_{i}\mid s_{i},a_{i};\bar{d}_{i},\theta)$ w/ RNF (Kobayashi and Aotani, 2023)
@@ 3.2 Adversarial learning with variational inference
First, adversarial learning defined in eq. (3) is redefined based on variational inference as a basis for incorporating the light robustness introduced in the next section. Specifically, we can focus on the fact that the disturbance $d$ is usually unobservable and regarded to be the latent variable. That is, by introducing the disturbance-aware model and the prior (a.k.a. nominal noise), the following evidence lower bound is derived according to Jensen’s inequality.
@@
where $\mathrm{KL}(\cdot||\cdot)$ denotes Kullback-Leibler divergence between two probabilities. When maximizing this lower bound, the adversary tries to minimize $\ln p_{w}(s^{\prime},r\mid s,a;d,\theta)$ to take into account the rare events, while regularizing $\varpi(d\ mids;\phi)\to\varpi(d)$. That is, the min-max problem in eq. (3) seems to be extended by adding the regularization.
The degree of regularization is commonly adjusted by introducing the gain $\beta\geq 0$ (Higgins et al., 2017) ($\beta=0$ means no regularization as like eq. (3)). Although this regularization is expected to reduce the generation of excessive disturbances, how it ($\beta$, more specifically) affects to the balance between robustness and conservativeness is unclear. For developing an auto-tuning mechanism of $\beta$ (or other alternative gain) to achieve the desired balance, some kind of criteria that can be intuitively specified by the user are needed.
In addition, we need to focus on the fact that the newly introduced disturbance-aware model, $p_{w}(s^{\prime},r\mid s,a;d,\theta)$, cannot be used when $d$ is unknown, so it is used only for adversarial learning. Only if an additional disturbance estimator could be introduced, it could be used even during inference to enable the domain adaptation. In that light, as the disturbance-marginalized model, $p_{w}(s^{\prime},r\mid s,a;\theta)$, is not included in the above lower bound, it should be additionally optimized by explicitly considering some kind of conditions.

### v2

@@ Abstract
journal: Robotics and Autonomous Systems
url]https://prinlab.org/en/
@@ 1 Introduction
Reinforcement learning (RL) attracts increasing attention as methodology for robots to learn stochastic control policies that enable them to accomplish a given task by trial and error [1]. In particular, model-based RL has high expectations for real-world robotic applications due to its excellent sample efficiency [2, 3], and various robotic applications have been reported recently: e.g. manipulation on complex dynamics of dexterous hand [4], cloth [5], and even grinding of object [6]; locomotion by quadruped [7], biped [8], and quadrotor [9]; and their combination [10]. Although the framework proposed in this study can be applied to model-free RL, for the sake of simplicity, this paper limits its focus to model-based RL.
@@
In addition to the sample efficiency (and, of course, control performance), recent RL studies have often aim to maximize robustness. This would enable robots to cope with unexpected situations in a faced environment, which can easily be caused by a variety of factors (e.g. biases in empirical data and unobservable information). Robustness is essential for robots to work in the real world “reliably” because it is difficult to observe all information there [11].
However, achieving robustness is not a free lunch, and in the context of RL, it often requires to experience a wide variety of situations, making its sample efficiency poor [12]. While the recent progresses have mitigate this problem by using the fast and parallel simulatior [24] and/or more flexible adversarial learning [27], the side effect of maximizing robustness, i.e. conservative policies that degrade control performance under normal conditions, is raised as a new problem [34]. Because this tradeoff is inherently unavoidable, another approach has been proposed in which, instead of increasing robustness, uncertainty is revealed online and reflected in models and/or policies [38]. In this case, the risk of erroneous uncertainty estimation is caused, and the computation cost during inference is increased due to estimating uncertainty, which affects real-time control. Although the details of these problems in related work are summarized in the next section, to my knowledge, no framework or theoretical system has yet been proposed that balances all of these issues.
In line with this background, this study proposes a new framework, so-called LiRA (Light-Robust Adversary) (see Fig. 1), for adversarial learning that satisfies the following two real-world requirements:
  1. 1.
  2. 2.
To this end, adversarial learning is first rederived according to variational inference [13]. Then, as a new definition of robustness, light robustness [14] is integrated with it. This relaxes full robustness by imposing an inequality constraint that limits the degradation from ideal performance due to disturbances within a specified threshold (this corresponds to the conservativeness). Through the Lagrangian method [15], this inequality constraint is transformed into a loss function for numerical optimization, allowing for automatic tuning of the adversary level. In addition, as such a learning process makes the model marginalized w.r.t. disturbances moderately robust, the disturbances can remain unknown at inference time, requiring no estimation cost or additional conditioning.
The behaviors of the proposed LiRA are confirmed by numerical simulations using Mujoco [16]. LiRA is not too conservative in control performance only with the nominal noise, and the degradation of control performance in response to disturbance intensity is suppressed. In addition, according to the specified conservativeness, the proposed LiRA attempts to increase the robustness by increasing the adversary level for the condition with weaker disturbance sensitivity, and vice versa. Such an auto-tuning capability also yields self-paced curriculum learning [17, 18], where disturbances are suppressed during the under-performance phase and reinforced as learning progresses. These results indicate that LiRA can achieve a good balance between robustness and conservativeness.
Finally, a real-robot learning using LiRA is demonstrated without simulations. By applying virtual external forces to its internal impedance model, a quadrupedal robot perceives disturbances only during learning. This task can be performed by LiRA only on the real robot with appropriate suppression of excessive disturbance, whereas a fully adversarial learning has a high risk of malfunctioning the robot. Although learning with nominal noise alone should not be sufficient to generate a force-responsive gait, LiRA makes it possible by moderately experiencing various situations during learning. Thus, this demonstration illustrates that LiRA can induce the robot to acquire the desired behaviors only with the limited data collected from the real world in short time (less than two hours in this demonstration).
@@
The contributions of this paper are threefold.
@@
    Theoretical aspect: By integrating the variational inference and light robustness, a novel learning rule, which provides the moderate robustness, is theoretically derived.
@@
    Implementation aspect: Three modules are introduced for the learning rule to function as expected, and their necessity is verified by simulations.
  3. 3.
    Practical aspect: By automatically adjusting the state-dependent adversary level, the efficient and safe adversarial learning is demonstrated in the real world, acquiring the robust policy.
Robustness of policies that do not impair control performance is critical in real-world applications, but its realization so far is based on the use of simulations that differ from reality in terms of efficiency and safety [19, 20]. On the other hand, this research, which guarantees efficiency and safety by using adversarial learning without excessive attacks, makes it possible to obtain robust policies even only in the real world. In other words, the proposed LiRA can be applied to problems that are difficult to simulate or take longer to simulate than in real time, and it can be adapted to real-world robotic systems that change in a non-stationary manner.
1.1 Related work
To enhance the robustness of RL policies, a naive approach is domain randomization on simulations [21, 22, 23], which encompasses a variety of optimal policies by learning from data experienced in simulations driven by various simulation parameters (friction, mass, texture, and so on). Although this approach has had much success in recent years [19, 20], it should be inefficient unless data are collected in parallel using plenty of computational resources, which waste huge amounts of energy [24, 25]. It is therefore difficult for real-world robots to continuously learn on-board with this approach. In addition, this approach implicitly assumes that the real world can be represented by either of the randomly selected simulation parameters [26], which requires either deep domain knowledge to find an appropriate parameter space or a broader parameter space to be tried. The former lacks generality, and the latter further increases the conservativeness of policies.
As a more efficient approach, an adversarial learning can be considered, in which an adversary is introduced that actively interferes with the robot’s task accomplishment [27, 28, 29]. Such active disturbances (modeled by highly expressive neural networks) can produce complex and challenging experiences more efficiently than random ones. However, this approach inevitably inherits the mode collapse issue in adversarial learning [30, 31], making learning the optimal policy unstable. In addition, in real-world robots, excessive disturbances might damage themselves and/or objects in the faced environment. The previous study that mitigates these issues is RRL-Stack [32], which is based on Stackelberg game and would be the most similar to the proposed method, but it requires additional rollouts with oracle policies, sacrificing the sample efficiency. Furthermore, one of the adverse effects of maximizing robustness (caused in both of the above) is to make the policy too conservative [33, 34, 35]. In other words, since robots have to take into account contingencies that may not occur in practice, they tend to choose safer and more secure actions. If the maximum disturbance intensity could appropriately be designed in advance, such conservativeness would be controlled at the minimum. However, as the relationship between the policy performance and the disturbance intensity is nonlinear and varies from situation to situation, this solution is infeasible.
This issue would be mitigated by domain adaptation [36, 37, 38], which estimates the uncertainty in the faced environment (e.g. represented as simulation parameters and disturbances) and incorporates them into the policy. As long as the uncertainty estimator is accurate, this approach can acquire the necessary and sufficient balance between robustness and conservativeness. However, its learning cost should be equal to or larger than that of domain randomization, since the effects of any potential uncertainty must be learned in advance. In addition, the extra computational cost of the uncertainty estimator and the uncertainty-conditioned policy can easily be an obstacle to real-world robot control, which is strictly time-constrained. In particular, it has been reported that in the model-based RL algorithms, where model predictive control (MPC) [3, 39] is employed in combination, reducing the extra computational cost as much as possible leads to better control performance because the MPC optimization converges in time [40, 41]. As the estimation error would also be raised as our concern, an alternative solution to this domain adaptation should be considered.
Thus, each of the conventional approaches has different drawbacks, as summarized in Table 1. Since all of these features are necessary for real-world applications, LiRA aims to achieve a good balance between all of them.
@@ 2.1 Model-based reinforcement learning
In RL [1], an agent aims to gain the maximum rewards from an environment in the future. For mathematical formulation of the relationship between the agent and environment, Markov decision process (MDP) is basically assumed, i.e. $(\mathcal{S},\mathcal{A},p_{e},r)$. Here, $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $p_{e}:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$ is the state transition probability (a.k.a. dynamics), and $r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathcal{R}$ is the reward function to evaluate each transition. Here, $\mathcal{R}\subseteq\mathbb{R}$ holds normally, but the case of multi-objective RL, where multiple rewards are simultaneously considered [42], is assumed for generality. That is, the output of reward function is vectorized as $\mathcal{R}\subseteq\ \mathbb{R}^{|\mathcal{R}|}$.
@@
If the agent knows $p_{e}$ and $r$ accurately, this problem can numerically be solved using MPC (in this paper, AccelMPPI [39] is employed). Therefore, model-based RL algorithms explicitly approximate them as a model ($p_{w}$ in this paper) using, for example, deep neural networks with parameters $\theta$ [2, 41].
@@ 2.2 Adversarial learning
To make $\pi^{\ast}$ obtained through the above optimization problem robust, worst-case scenarios for the objective function (i.e. the sum of predicted reward, called the return) are basically considered [43, 44]. However, this incurs an extra computational cost at inference time compared to the case simply using expected value of the prediction. In addition, using these methods, it is difficult to determine whether robustness is due to the model or to such a robust MPC. This study therefore focuses on the direction that robust control is achieved by making $p_{w}$ robust during learning, referring to the result reported in the literature [45] that control performance becomes implicitly robust if $p_{w}$ is optimized in consideration of events that are rare in reality.
@@ 3.1 Overview
The basic adversarial learning suffers from the mode collapse issue [30, 31], where the min-max problem becomes unbalanced (in most cases, the adversary becomes dominant), preventing the model from robust learning properly. In order to improve the robustness of the control performance to the extent that it does not become too conservative, this paper proposes LiRA. LiRA utilizes two theoretical tricks in its formulation.
First, the adversarial learning in eq. (3) is reformulated by variational inference [13], introducing the necessary components in a natural way. In particular, the function to regularize the adversary to its prior distribution is derived, thus inhibiting excessive attacks by the adversary, albeit passively. Then, the light robustness [14] is applied as an additional constraint for the gap of predictive performance of models with/without disturbance information, which can be converted to a regularization term according to Lagrangian method with an auto-tuned weight parameter. This parameter has a role of adjusting adversary level automatically, actively balancing the robustness and conservativeness.
To appropriately solve the derived optimization problem, this paper implements three practical tricks (see the section 4). First, noting that the bias due to disturbances should ideally be zero, a design constraint is provided by a restricted normalizing flows (RNF) [46] such that the expected values of the two models are consistent with each other. Second, to make a gradient reversal layer (GRL) [47] available for the stable and efficient adversarial learning, a trick to hindsight the computational graph from the disturbance generator to the model via the disturbance, so-called hindsight reparameterization gradient (HRG), is newly developed. Third, a new equality constraint to actively reflect the worst cases for either the model or the disturbance generator, midrange-mean balancing (MMB), is given, following the literature [45], which efficiently improves the robustness of the model by (moderately) prioritizing the worst-case data for the performance.
The details of the above proposal will be presented in the following sections. Especially, its process for computing the loss function to be minimized is summarized in Fig. 2. The overall pseudocode of LiRA is also shown in Alg. 1. Note that the loss to be minimized in this pseudocode, $\mathcal{L}^{x}$, has the computational graph only for $x$ and cuts for the other variables to be optimized. In addition, $\mathcal{L}^{\phi}$ corresponds to Kullback-Leibler divergence approximated by Monte Carlo method, which is necessary as the adversary is designed by normalizing flows.
Figure 2: Process to calculate loss of LiRA: the dashed lines show the non-backpropagatable signals replayed from the buffer; the blue lines and blocks show per-sample processing, while the green lines and blocks are for processing for the batch, making values of samples a scalar loss.
@@
19:         $\ell^{a}_{i}=-\ln p_{w}(s^{\prime}_{i},r_{i}\mid s_{i},a_{i};\bar{d}_{i},\theta)$ w/ RNF [46]
@@ 3.2 Adversarial learning with variational inference
First, adversarial learning defined in eq. (3) is redefined based on variational inference as a basis for incorporating the light robustness introduced in the next section. Specifically, we can focus on the fact that the disturbance $d$ is usually unobservable and regarded to be the latent variable. In maximizing the log-likelihood of disturbance-marginalized model, $p_{w}(s^{\prime},r\mid s,a;\theta)$, in which the unobservable $d$ is excluded by marginalization, maximizing its worst case is more appropriate from the point of view of robustness. That is, by introducing the disturbance-aware model and the prior (a.k.a. nominal noise), the following evidence lower bound is derived according to Jensen’s inequality (and inequality between expected value and infimum).
@@
where $\mathrm{KL}(\cdot||\cdot)$ denotes Kullback-Leibler divergence between two probabilities. When maximizing this lower bound, the adversary tries to minimize $\ln p_{w}(s^{\prime},r\mid s,a;d,\theta)$ to take into account the rare events, while regularizing $\varpi(d\ mids;\phi)\to\varpi(d)$. That is, the min-max problem in eq. (3) seems to be extended by adding the regularization, which might alleviate the mode collapse with the excessive attacks by the adversary.
The degree of regularization is commonly adjusted by introducing the gain $\beta\geq 0$ [48] ($\beta=0$ means no regularization as like eq. (3)). Although this regularization is expected to reduce the generation of excessive disturbances, how it ($\beta$, more specifically) affects to the balance between robustness and conservativeness is unclear. For developing an auto-tuning mechanism of $\beta$ (or other alternative gain) to achieve the desired balance, some kind of criteria that can be intuitively specified by the user are needed.
In addition, we need to focus on the fact that the newly introduced disturbance-aware model, $p_{w}(s^{\prime},r\mid s,a;d,\theta)$, cannot be used when $d$ is unknown, so it is used only for adversarial learning. Only if an additional disturbance estimator could be introduced, it could be used even during inference to enable the domain adaptation. In that light, as $p_{w}(s^{\prime},r\mid s,a;\theta)$ is not included in the above lower bound, it should be additionally optimized by explicitly considering some kind of conditions.